Suppose the following information was taken from the 2017 financial statements of FedEx Corporation, a major global transportation/delivery company.

(in millions)

2017

2016

Accounts receivable (gross)

$ 3,704

$ 4,562

Accounts receivable (net)

3,263

4,069

Allowance for doubtful accounts

441

493

Sales revenue

37,499

35,432

Total current assets

7,754

7,016


Answer each of the following questions.

Calculate the accounts receivable turnover and the average collection period for 2017 for FedEx Corporation. (Round answers to 1 decimal place, e.g. 12.5. Use 365 days for calculation.)

Accounts receivable turnover times
The average collection period for 2017 days

eTextbook and Media

List of Accounts

Is accounts receivable a material component of the company’s total current assets? Choose your answer here                                                                      YesNo

eTextbook and Media

List of Accounts
Answer #1
Accounts receivable turnover = Sales revenue/Average Accounts receivable (net)
Accounts receivable turnover 10.2 times =37499/((3263+4069)/2)
The average collection period for 2017 35.8 or 35.7 days =365/10.2
2
% of Accounts Receivable to Current Assets 42% =3263/7754
Accounts receivable are a signification portion of Total current assets
Yes, accounts receivable is a material component of the company’s total current assets